A single-phase load takes 55 kw at 70% pf lagging from a 240V, 50 HZ supply. If the supply is made 60 Hz, with the voltage remaining the same what will be the KW load at 60 Hz?
1
Expert's answer
2021-02-25T04:08:38-0500

Explanation & Calculations


  • Since there is a power factor given ("\small \cos\theta =0.7\,\text{lagging}") it can be understood that the load is not purely resistive, rather some inductance & capacitance is present.
  • Therefore, the behavior of the load is frequency dependant.
  • To calculate the new power the new current is needed but there is not enough information to estimate the reactance involved in the load.
  • All we can do is generating information from the basic information given: AC voltage, lagging current.
  • If we write for the instantaneous voltage, current & power,

"\qquad\qquad
\begin{aligned}
\small V&=V_m\sin\omega t\qquad i=i_m\sin(\omega t-\theta)\\
\small P&= Vi\\
&= \small V_m\sin\omega t\cdot i_m\sin(\omega t-\theta)\\
&= \small \frac{V_mi_m}{2}\cdot\big[cos\theta-\cos(2\omega t-\theta)\big]\\
&= \small V_{rms}i_{rms}\cos\theta-V_{rms}i_{rms}\cos(2\omega t-\theta)
\end{aligned}"


  • As it is seen the instantaneous power consists of two components: one is frequency-independent & the other is frequency-dependent.
  • Since it is the average power that is considered long term wise, the behavior of "\small \cos(2\omega t-\theta)" counts to zero yielding the power of the circuit to be

"\qquad\qquad
\begin{aligned}
\small P&= \small Vi\cos \theta
\end{aligned}"

which remains constant regardless of the frequency.

  • It is the active power on the other hand when the average is considered.


  • What is asked in the question is the active power (as it is given in Watts).
  • Therefore, the average\active power does not change even when the frequency is changed to 60hz.
  • It is 55 kW.